The point (-3/5, y) in the third quadrant corresponds to angle θ on the unit circle.
The value of sec θ is ? , and the value of cot θ is ?
so what are the values of secθ and cotθ?

Answer by stanbon(75887)

The point (-3/5, y) in the third quadrant corresponds to angle θ on the unit circle.
-------------------
r = sqrt[y^2-(3/5)^2] = sqrt[(25y^2-9)/25]
--------------------------------------------
The value of sec θ is ?
sec(t) = x/r = (-3/5)/sqrt[(25y^2-9)/25]
===========================================
the value of cot θ is ?
cot(t) = x/y = (-3/5)/y = -3/(5y)
--------------------------------------------
Cheers,
Stan H.
--------------